the difference between the measurements of two complementary angles is 18 degrees what is the measurement of each angles
Answer by algebrahouse.com(1659)

"the difference between the measurements of two complementary angles is 18 degrees what is the measurement of each angles"

x = one angle
90 - x = other angle {complimentary angles add up to 90}

90 - x - x = 18 {difference between the two angles is 18}
90 - 2x = 18 {combined like terms}
-2x = -72 {subtracted 90 from both sides}
x = 36 {divided both sides by -2}
90 - x = 54 {substituted 36 in for x, into 90 - x}

54 and 36 are the two angles
www.algebrahouse.com